### v1

# Effects of retardation on many-body superradiance in chiral waveguide QED

Abstract
We study the superradiant decay of a chain of atoms coupled to a chiral waveguide, focusing on the regime of non-negligible photon propagation time. Using an exact master equation description which accounts for delay effects, we obtain evidence to suggest that competition between collective decay and retardation leads to the emergence of an effective maximum number of atoms able to contribute to the superradiant dynamics, resulting in a plateau of the peak emission rate. To develop this analysis further, we investigate the inter-atomic correlations to find features consistent with the formation of individual superradiant domains. Moreover, we find that retardation can also result in persistent oscillatory atomic dynamics accompanied by a periodic sequence of emission bursts.
A major objective of modern quantum optics is to characterise cooperative phenomena, one of the most striking examples being superradiance [1], the emission of a short radiation burst from a collection of $N$ emitters, with a peak intensity scaling superlinearly in $N$. Since the first formulation of the theory of superradiance [2], there has been significant interest in understanding superradiant dynamics in extended media [3, 4, 5, 6, 7, 8], where the build-up of coherence among atoms may be undermined by their spatial separation. In particular, a key question which remains open is how superradiance is affected in the regime where propagation times of the photons mediating dissipative interactions are non-negligible, rendering the dynamics non-Markovian.
A novel setting for exploring superradiance is that of waveguide quantum electrodynamics (WQED). Here, long-range dissipative interactions mediated by photonic modes confined in a waveguide enable the superradiant decay of distant emitters [9, 10, 11], signatures of which have already been observed in a variety of experimental platforms [12, 13, 14, 15, 16, 17, 18]. In suitably structured waveguides, the group velocity of the guided modes can be engineered [19, 20] to enter the regime of non-negligible retardation in a controlled manner. Some theoretical works have already explored the non-Markovian effects which arise in this slow-light WQED setting [21, 22, 23, 24, 25, 26], however these studies were limited to a few excitations, since simulating many-body dynamics beyond the Markovian approximation is generally exceedingly difficult.
As we will demonstrate, this is not true in the case of a chiral waveguide, where the photon-mediated interactions are unidirectional [27, 28, 29]. Here, the emitter-waveguide setup constitutes a cascaded quantum system [30, 31, 32], which is much more tractable. Recently, a superradiant burst from $N\approx 10^{3}$ atoms in such a system was experimentally observed [17, 33], and chiral WQED therefore presents a unique vantage point for exploring the effects of retardation on superradiance, combining the possibility of rigorous theoretical investigation with a proximity to experiments.
In this Letter, we show that non-negligible photon propagation times can significantly alter the collective decay dynamics in a chiral WQED setting. Our simulations, based on an exact master equation description, reveal that retardation suppresses the characteristic superradiant scaling of the peak emission rate, leading instead to a plateau in peak emission along the chain of waveguide-coupled atoms. This signals the emergence of a maximum cooperative system size, resulting in local rather than global synchronisation among the atoms. We characterise this further by studying the atomic correlations and supplement our numerical observations by mean-field analyses. Finally, we demonstrate that sufficiently long chains support sustained oscillatory dynamics, resulting in a periodic emission of intensity bursts.
Model & exact master equation
We consider a one-dimensional linear waveguide and $N$ two-level atoms coupled chirally to the right-propagating guided modes, described by the Hamiltonian ($\hbar=1$[34, 35, 36, 27, 37]
[…]
where $v>0$ is the group velocity of the guided modes, $\Gamma$ is the single-atom decay rate into the waveguide, and $x_{n}$ denotes the position of the $n$th atom. We label the atoms according to their position, with $x_{n}<x_{n+1}$.
Moving beyond the Markovian assumption of negligible photon propagation times typically requires explicit simulation of the environment. However, in cascaded settings such as the one considered here, the absence of feedback from atoms further downstream allows us to trace out the waveguide exactly. To this end, we introduce a time-shifted picture, where the time axis for atomic operators is shifted depending on the atomic position. Explicitly, for an arbitrary operator $O_{n}$ acting on the $n$th emitter, the relation between the Heisenberg picture and the time-shifted picture (denoted with a hat) is $\hat{O}_{n}(t)=O_{n}(t-t_{n})$, where $t_{n}=(x_{N}-x_{n})/v$. Alternatively, we can define the atomic density operator in the time-shifted picture, $\hat{\rho}(t)$, which evolves as (see Supplementary Material (SM) [38]),
$\begin{split}\partial_{t}\hat{\rho}(t)&=\sum_{n}\Theta(t-t_{n})\mathcal{L}_{n}\hat{\rho}(t)\\ \mathcal{L}_{n}\hat{\rho}(t)&=\frac{\Gamma}{2}[\sigma_{n}\hat{\rho}(t),\sigma^{\dagger}_{n}]+\Gamma\sum_{m>n}[\sigma_{n}\hat{\rho}(t),\sigma^{\dagger}_{m}]+\mathrm{H.c.}\end{split}$ (2)
Although it is well established that delay times can be accounted for through suitable frame transformations [32], the physical consequences of this have hardly been explored. In particular, most previous works have focused on steady-state properties or sufficiently long-time dynamics [23, 39, 31, 30, 40], for which time delays have no significant impact. However, in the context of superradiance, they play a crucial role.
In the following, we assume the emitters to be equidistantly arranged with spacing $d$, and denote by $\tau=d/v$ the time taken for photons to propagate between neighbouring atoms. We focus on initial states with no photons in the waveguide and the emitters in a pure state $\rho(t=0)=\ket{\psi_{0}}\!\bra{\psi_{0}}$ |, where
|$\ket{\psi_{0}}=\left[\cos\left(\frac{\theta_{0}}{2}\right)\ket{g}+\sin\left(\frac{\theta_{0}}{2}\right)\ket{e}\right]^{\otimes N}.$ (3)
We study the collective decay of these states, after the atom-photon coupling is turned on at time $t=0$, as accounted for by the heaviside step function $\Theta(t)$ in Eq. (1). Specifically, we compute local emission rates $r_{n}(t)=-\partial_{t}\langle\sigma_{n}^{\dagger}\sigma_{n}\rangle_{t}$, with $\langle\ldots\rangle_{t}\equiv\mathrm{tr}(\ldots\rho(t))$, noting that the dynamics of any emitter is independent of the presence of more emitters further downstream, so $r_{n}(t)$ is insensitive to the total number of atoms $N$.
Even with the emitter-only description of the exact dynamics through Eq. (2), its simulation still presents a considerable computational challenge. In this work, we tackle this problem using simulations based on the Truncated Wigner Approximation (TWA) [41, 42, 43, 44, 45, 46] (see SM [38] for details). In the large-$N$ limit, Dicke superradiance is known to consist of classical evolution driven by initial quantum fluctuations [47], which aligns ideally with the approximations of the TWA [48, 49, 50, 51, 52, 53, 54]. Since chiral interactions and time delays break the permutation symmetry of the Dicke problem, we further establish the accuracy of our TWA simulations by benchmarking against quantum trajectory simulations with Matrix Product States (MPS), finding excellent agreement (see SM [38]).
Figure 1: Effect of retardation for a partially-inverted state with $\theta_{0}=0.7\pi$, computed using TWA (markers) and MFT (solid lines). (a) Local emission rate $r_{n}(t)$ at $n=200$, evaluated for various time delays $\tau$. (b) Scaling of $r^{\mathrm{pk}}_{n}$ with $n$ for various $\tau$, with the values of $\tau$ shown in panel (a) indicated by the corresponding colours. (c) Comparison of the peak emission rate $r_{\mathrm{eff}}$ and time $t_{\mathrm{eff}}$ with the analytical predictions (4).
Figure 2: (a)-(c) Same as Figs. 1a-c for a fully-inverted initial state, with $n=150$ in (a). (d) Time-evolution of the correlator $C(n,n-j,t)$ at $n=150$, for $\tau=0$ (top) and $\Gamma\tau=10^{-3}$ (bottom). In both panels, we overlay $r_{n}(t)$ for qualitative reference and also indicate the peak emission time $t^{\mathrm{pk}}_{n}$. In the lower panel, we also indicate the light cone. (e) Scaling of $\xi^{\mathrm{pk}}_{n}$ with $n$ for increasing $\tau$, with the values of $\tau$ shown in panel (a) indicated by the corresponding colours. (f) Scaling of the asymptotic correlation length $\xi_{\mathrm{eff}}$ with $\tau$, with the small-$\tau$ power law scaling indicated by the dashed line.
Partially-inverted initial states
In Fig. 1a, we show emission dynamics for a partially-inverted initial state ($\theta_{0}<\pi$ in Eq. (3)) for different time delays $\tau$. We observe the characteristic signature of a superradiant burst, namely a peak value $r_{n}^{\mathrm{pk}}$ of the site-$n$ emission rate $r_{n}(t)$ exceeding the single-atom maximum value, $r^{\mathrm{pk}}_{n}>\Gamma$, at a delayed peak emission time $t=t^{\mathrm{pk}}_{n}>0$. With respect to the retardation-free case, the effect of the retardation is to delay the peak emission further, resulting in a superradiant burst with a lower amplitude $r^{\mathrm{pk}}_{n}$ at a later time $t^{\mathrm{pk}}_{n}$. This reflects the fact that a non-negligible retardation time allows for atoms downstream to decay independently before the light emitted from previous atoms impinges on them, reducing the buildup of the superradiant pulse along the waveguide. Note also that $r_{n}(t=0)>\Gamma$ for $\tau=0$, since a state (3) with $\theta_{0}<\pi$ contains (classical) correlations between atoms, however for $\tau>0$, their effect can only manifest progressively.
In Fig. 1b, we show the scaling of the peak emission rate $r^{\mathrm{pk}}_{n}$ with $n$. In the retardation-free case, we observe the Dicke scaling $r^{\mathrm{pk}}_{n}\sim n\Gamma$ [2, 1], while for $\tau>0$, the value of $r^{\mathrm{pk}}_{n}$ instead plateaus to a maximum effective emission rate $r_{\mathrm{eff}}$. A similar effect can be seen for the peak emission time $t^{\mathrm{pk}}_{n}$, which plateaus to some $t_{\mathrm{eff}}$. This suggests that the competition between retardation and cooperative emission leads to an effective maximum number of atoms $N_{\mathrm{eff}}$ which can contribute to the superradiant decay. This maximum cooperative system size can be estimated by a self-consistency argument: the time taken for photons emitted by the first atom to reach the $n$-th atom along the waveguide is $n\tau$, while this atom decays superradiantly on a timescale $\sim 1/(n\Gamma)$. Equating these times, the effective number of atoms upstream from the $n$th atom which can enhance its emission should scale as $N_{\mathrm{eff}}\sim(\Gamma\tau)^{-1/2}$, such that the Dicke relations $r_{n}^{\mathrm{pk}}\sim n\Gamma$ and $t_{n}^{\mathrm{pk}}\sim 1/(n\Gamma)$ imply
$r_{\mathrm{eff}}\sim(\tau/\Gamma)^{-1/2}\quad\text{and}\quad t_{\mathrm{eff}}\sim(\tau/\Gamma)^{1/2}\,.$ (4)
While these scalings were already proposed in Ref. [4], we are now in a position to confirm them numerically. In Fig. 1c, we show the dependence of $r_{\mathrm{eff}}$ and $t_{\mathrm{eff}}$ on $\tau$, and find that, indeed, at sufficiently small $\Gamma\tau$, the scalings (4) are observed. At larger delays, however, we find that for sufficiently large initial inversion angles $\theta_{0}$, the scalings diverge from the prediction (4). This can be understood by noting that the scalings $r_{n}^{\mathrm{pk}}\sim n\Gamma$ and $t_{n}^{\mathrm{pk}}\sim 1/(n\Gamma)$ hold for $n\gg 1$, however at large $\Gamma\tau$ the effective cooperative system size $N_{\mathrm{eff}}$ becomes small enough for these scalings to break down. This effect is more pronounced for larger $\theta_{0}$, where the asymptotic scalings are approached more slowly.
Mean-field analysis
For partially-inverted initial states, the emission dynamics are also captured by a mean field theory (MFT) description in the limit of small delay times ($\Gamma\tau\ll 1$), where the atomic array can be treated as a continuous medium [3, 6, 4, 52, 7, 8, 1]. We express $\langle\sigma_{n}\rangle_{t}=\frac{1}{2}\sin\theta(n,\Gamma t)$ and $\langle\sigma_{n}^{z}\rangle_{t}=-\frac{1}{2}\cos\theta(n,\Gamma t)$, in terms of a continuous bivariate function $\theta(x,t)$ with dimensionless arguments obeying (see SM [38] and Refs. [55, 56])
$\left[\left(\Gamma\tau\right)\frac{\partial^{2}}{\partial t^{2}}+\frac{\partial^{2}}{\partial t\partial x}\right]\theta(x,t)+\sin\theta(x,t)=0\,,$ (5)
with boundary conditions $\theta(x,0)=\theta(0,t)=\theta_{0}$. A numerical solution using the method of finite differences [57] shows excellent agreement of the emission dynamics with the TWA numerics for sufficiently small delays (see Figs. 1a,b).
Moreover, Eq. (5) itself serves as a starting point for some insightful scaling arguments: For $\tau=0$, Eq. (5) becomes the Sine-Gordon equation [58] $\partial_{t}\partial_{x}\theta(x,t)+\sin\theta(x,t)=0$, which is symmetric under $x\leftrightarrow t$. With our boundary conditions, it can be solved as $\theta(x,t)=u(xt)$ with $z\,u^{\prime\prime}(z)+u^{\prime}(z)+\sin u(z)=0$, subject to the initial conditions $u(z=0)=\theta_{0}$ and $u^{\prime}(z=0)=-\sin\theta_{0}$ [4]. Note that this change of variables implies $\theta_{n}(t)=\theta_{m}(nt/m)$ and hence $r_{n}(t)=(n/m)\,r_{m}(nt/m)$, which is a statement of the scalings for the retardation-free case assumed above. More generally, we can see from Eq. (5) that the solution $\theta(x,t;\tau)$ for a given delay $\tau$ is related to the solution for a delay $\alpha\tau$ ($\alpha>0$) as $\theta(x,t;\alpha\tau)=\theta(\sqrt{\alpha}x,t/\sqrt{\alpha};\tau)$. Defining the associated emission rate $r(x,t;\tau)=-\frac{1}{2}\sin\theta(x,t;\tau)\partial_{t}\theta(x,t;\tau)$, we can then see that $r(x,t;\alpha\tau)=r(\sqrt{\alpha}x,t/\sqrt{\alpha};\tau)/\sqrt{\alpha}$. From these relations, it follows directly that the asymptotic scalings in Eq. (4) are exact at the mean-field level (see Fig. 1c) and the departure from these scalings at larger $\Gamma\tau$ is consistent with the breakdown of the continuum approximation in this regime.
Fully-inverted initial state
Beyond capturing departures from mean-field predictions, our numerics also allow us to study the fully-inverted initial state |$\ket{\psi_{0}}=\ket{e}^{\otimes N}$ conventionally considered in Dicke superradiance, which is adynamical under Eq. (5), since the decay of this state is initialised by vacuum fluctuations not accounted for by MFT [47]. In Figs. 2a,b we see that the plateau effect described above for partially-inverted states manifests analogously in this case. However, the scaling of $t_{\mathrm{eff}}$ displays a significant divergence from the predicted scaling. This can be understood by noting that the correct $n$-dependence of $t^{\mathrm{pk}}_{n}$ in the retardation-free case is actually given by $t^{\mathrm{pk}}_{n}=c_{1}\log(c_{2}n)/n$, with some constants $c_{1},c_{2}$ [59]. Indeed, we find that the scaling $t_{\mathrm{eff}}=c_{3}(\tau/\Gamma)^{1/2}\log(c_{4}/(\Gamma\tau))$ implied by this is a much better fit (see Fig. 2c).
We now consider the implications of the emergent maximum cooperative system size $N_{\mathrm{eff}}$ for a chain of $N\gg N_{\mathrm{eff}}$ atoms: intuitively, such a chain can be expected to “split” into individual regions of size $\sim N_{\mathrm{eff}}$, each decaying superradiantly. As a starting point for characterising this domain structure, we introduce the multi-time correlator
$C(n,m,t)=\langle\sigma_{n}^{\dagger}(t)\sigma_{m}(t-(n-m)\tau)\rangle_{0}\,,$ (6)
which is real-valued for the fully-inverted initial state. In Fig. 2d, we show the time-evolution of $C(n,n-j,t)$. In the retardation-free case, we observe a buildup of correlations around $t\approx t^{\mathrm{pk}}_{n}$, reflecting the emergent coherence across the atoms known to accompany superradiant decay [1, 2]. Contrary to the Dicke case, however, these correlations are not uniform, but decrease with distance $j$ from the $n$th site.
The case of $\tau>0$ displays two important differences with respect to the retardation-free case: Firstly, Fig. 2d shows the emergence of a “light cone” structure, since $C(n,n-j,t)=0$ for $t<j\tau$. Secondly, the correlations at peak emission decay with $j$ over shorter length scales along the atomic array. To make this second observation more quantitative, we note that for $t\leq t^{\mathrm{pk}}_{n}$, the correlations are well-captured by a compressed exponential fit $C(n,n-j,t)=C_{0}e^{-(jd/\xi_{n}(t))^{\alpha}}$ (see SM [38]), where the fit parameter $\xi_{n}(t)$ can be interpreted as a (time-dependent) correlation length. In the retardation-free case, we find that its value at $t=t^{\mathrm{pk}}_{n}$, which we denote by $\xi^{\mathrm{pk}}_{n}=\xi_{n}(t=t_{n}^{\mathrm{pk}})$, scales as $\xi^{\mathrm{pk}}_{n}\sim nd$. For $\tau>0$, we can see that $\xi^{\mathrm{pk}}_{n}$ instead approaches a fixed asymptotic value $\xi_{\mathrm{eff}}$ just like $r^{\mathrm{pk}}_{n}$ and $t^{\mathrm{pk}}_{n}$ (see Fig. 2e). This asymptotic $n$-independent correlation length provides a direct signature of the formation of superradiant domains: it implies that in a large system, any atom sufficiently far along the chain can synchronise only across a finite range $\sim\xi_{\mathrm{eff}}$, leading to local rather than global collective decay. Additionally, we find that $\xi_{\mathrm{eff}}$ displays the scaling $\xi_{\mathrm{eff}}/d\sim(\Gamma\tau)^{-1/2}$ at sufficiently small $\tau$ (see Fig. 2f), in accordance with our interpretation of a maximum cooperative system size.
Multi-peak structure
Finally, while in the preceding discussion we have focused entirely on the properties of the first maximum of $r_{n}(t)$, retardation also leads to intriguing novel features in the emission dynamics at later times. For $\tau>0$, our preceding discussion has shown that the evolution of all atoms sufficiently far down the chain will be determined by an identical environment of a fixed number of atoms further upstream. At the level of the MFT, this can be formalised as $\theta(x\to\infty,t)=\theta_{\infty}(t)$, where $\theta_{\infty}(t)$ denotes the solution to the equation obtained by dropping the spatial derivative in Eq. (5), i.e. $(\Gamma\tau)\ddot{\theta}_{\infty}(t)+\sin\theta_{\infty}(t)=0$. The solution to this equation can be obtained analytically (see SM [38]) and displays a periodic time-dependence: $\theta_{\infty}(t)$ oscillates between ±$\pm\theta_{0}$, while the asymptotic (uniform) local emission rate $r(x\to\infty,t)=-\frac{1}{2}\sin\theta_{\infty}(t)\dot{\theta}_{\infty}(t)\equiv r_{\infty}(t)$ oscillates between ±$\pm r_{\mathrm{eff}}$. In Fig. 3, we compare this mean-field prediction against our numerics by examining the site-$n$ emission rate $r_{n}(t)$ for $n\gg 1$. We observe that indeed $r_{n}(t)\approx r_{\infty}(t)$ until a time $t\sim n\tau$, at which point the finite-site effects propagating from the beginning of the emitter chain reach the $n$th site. A similar $n$-independent periodic behaviour can be seen for the field intensity in the waveguide immediately after the $n$th site, $I_{n}(t)=\langle b^{\dagger}(x_{n}^{+})b(x_{n}^{+})\rangle_{t}$ (with $x^{+}\equiv x+0^{+}$), which can be expressed in terms of local emission rates as
$I_{n}(t)=\frac{1}{v}\sum_{m\leq n}\Theta(t-(n-m)\tau)\,r_{m}(t-(n-m)\tau)\,.$ (7)
Specifically, the oscillations in $r_{n}(t)$ are accompanied by a sequence of emission bursts in $I_{n}(t)\approx I_{\infty}(t)$ (see Fig. 3).
The prediction of infinitely sustained periodic dynamics in the thermodynamic limit raises intriguing questions. Firstly, the precise physical mechanisms underlying the phenomenon remain to be understood. Secondly, the fate of the oscillations beyond MFT is unclear, since the errors due to neglected correlations at the mean-field level accumulate over time, and in a many-body system like the one considered here, thermalisation effects could be expected to wash out coherent dynamics at late times. We leave a detailed investigation of these questions to future work, and emphasise here only that the dynamics in Fig. 3 represent a departure from the well-studied retardation-free superradiant dynamics.
Figure 3: Dynamics of the site-$n$ emission rate $r_{n}(t)$ (top row) and local field intensity $I_{n}(t)$ (bottom row) for $\theta_{0}=\pi/2$ and $\Gamma\tau=5\times 10^{-4}$, evaluated at various $n$ using TWA (solid blue lines). We also indicate the asymptotic mean-field predictions (dashed black lines) and the time $t=n\tau$ (dashed gray lines).
Conclusion
In this Letter, we have shown that non-negligible photon propagation times can lead to striking qualitative departures from the standard theory of superradiance, by inhibiting the global synchronisation of the atoms typically associated with superradiant decay. An effective maximum cooperative system size emerges, which manifests in a suppression of the characteristic superlinear scaling of the peak emission rate with system size, and in the formation of locally synchronised superradiating domains. Building on a mean-field description, we have also identified regimes in which retardation induces sustained periodic atomic dynamics which manifest in a regular sequence of emission bursts.
The contribution of our work is two-fold: Firstly, we have provided a partial answer to the long-standing problem of superradiance in extended atomic media, relying on sophisticated numerical simulations beyond qualitative conjectures. In doing so, we have also identified the setting of slow-light chiral WQED as a potential experimental platform for exploring this regime in a controlled manner. Secondly, our results provide a first glimpse into the varied phenomena that could be observed in collective decay dynamics beyond the Markov approximation. Suitably structured photonic environments can give rise to a host of non-Markovian phenomena, which have only been explored in the single-excitation regime [60, 61, 62, 63, 64, 65], and our work is a timely addition to the nascent research effort to explore the many-body dynamics of such systems.
Acknowledgements
We are very grateful to C. Mink for generous advice on the implementation of the TWA numerics. B.W., M.B., and J.I.C. acknowledge funding from the Munich Center for Quantum Science and Technology (MCQST), funded by the Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy (EXC2111-390814868). D. M. acknowledges support from the Novo Nordisk Foundation under grant numbers NNF22OC0071934 and NNF20OC0059939.
Note added
During the completion of this manuscript, a related work applying the TWA to superradiant decay in chiral WQED appeared in Ref. [66], which however focuses exclusively on the retardation-free case.

### v2

@@
 We label the atoms according to their position, with $x_{n}<x_{n+1}$.
+We assume here perfect coupling into the waveguide but demonstrate the robustness of our results in the presence of additional decay into unguided modes in the End Matter.
 Moving beyond the Markovian assumption of negligible photon propagation times typically requires explicit simulation of the environment.
@@
 In Fig.
-1c, we show the dependence of $r_{\mathrm{eff}}$ and $t_{\mathrm{eff}}$ on $\tau$, and find that, indeed, at sufficiently small $\Gamma\tau$, the scalings (4) are observed.
-At larger delays, however, we find that for sufficiently large initial inversion angles $\theta_{0}$, the scalings diverge from the prediction (4).
+1c, we show the scalings of $r_{\mathrm{eff}}$ and $t_{\mathrm{eff}}$ with $\tau$, and find that, indeed, at sufficiently small $\Gamma\tau$, the scalings (4) are observed.
+At larger delays, however, we find that for sufficiently large initial inversion angles $\theta_{0}$, the scalings diverge from the predicted scalings.
 This can be understood by noting that the scalings $r_{n}^{\mathrm{pk}}\sim n\Gamma$ and $t_{n}^{\mathrm{pk}}\sim 1/(n\Gamma)$ hold for $n\gg 1$, however at large $\Gamma\tau$ the effective cooperative system size $N_{\mathrm{eff}}$ becomes small enough for these scalings to break down.
@@
 Mean-field analysis
-For partially-inverted initial states, the emission dynamics are also captured by a mean field theory (MFT) description in the limit of small delay times ($\Gamma\tau\ll 1$), where the atomic array can be treated as a continuous medium [3, 6, 4, 52, 7, 8, 1].
+For partially-inverted initial states, the emission dynamics are also captured by a mean field theory (MFT) description in the limit of small delay times ($\Gamma\tau\ll 1$), where the atomic array can be treated as a continuous medium [3, 6, 4, 52, 7, 8, 1, 55, 56].
 We express $\langle\sigma_{n}\rangle_{t}=\frac{1}{2}\sin\theta(n,\Gamma t)$ and $\langle\sigma_{n}^{z}\rangle_{t}=-\frac{1}{2}\cos\theta(n,\Gamma t)$, in terms of a continuous bivariate function $\theta(x,t)$ with dimensionless arguments obeying (see SM [38] and Refs.
@@
 However, the scaling of $t_{\mathrm{eff}}$ displays a significant divergence from the predicted scaling.
-This can be understood by noting that the correct $n$-dependence of $t^{\mathrm{pk}}_{n}$ in the retardation-free case is actually given by $t^{\mathrm{pk}}_{n}=c_{1}\log(c_{2}n)/n$, with some constants $c_{1},c_{2}$ [59].
+This can be understood by noting that the correct $n$-dependence of $t^{\mathrm{pk}}_{n}$ in the retardation-free case is actually given by $t^{\mathrm{pk}}_{n}=c_{1}\log(c_{2}n)/(n\Gamma)$, with some constants $c_{1},c_{2}$ [61].
 Indeed, we find that the scaling $t_{\mathrm{eff}}=c_{3}(\tau/\Gamma)^{1/2}\log(c_{4}/(\Gamma\tau))$ implied by this is a much better fit (see Fig.
@@
 During the completion of this manuscript, a related work applying the TWA to superradiant decay in chiral WQED appeared in Ref. [68], which however focuses exclusively on the retardation-free case.
+END MATTER
+In realistic settings, the waveguide-coupled atoms are typically also subject to independent decay into unguided free-space modes, which competes with the correlated emission into the waveguide.
+We account for this effect by supplementing the unitary dynamics under the Hamiltonian (1) with a local dissipation term $\gamma\sum_{n}\mathcal{D}[\sigma_{n}]$ with $\mathcal{D}[x]~{}\bm{\cdot}=x\bm{\cdot}x^{\dagger}-\{x^{\dagger}x,\bm{\cdot}\}/2$.
+At the level of the exact master equation (2), this amounts to the modification $\mathcal{L}_{n}\hat{\rho}(t)\to\mathcal{L}_{n}\hat{\rho}(t)+\gamma\mathcal{D}[\sigma_{n}]\hat{\rho}(t)$.
+Here, we discuss the effect of this additional loss mechanism on the retardation-induced peak emission plateau.
+Specifically, we consider local emission rates into the waveguide, which we define as $r_{n}^{\mathrm{wg}}(t)=r_{n}(t)-\gamma\langle\sigma_{n}^{\dagger}\sigma_{n}\rangle_{t}$.
+In the absence of retardation, superradiant decay of the $n$-th atom (i.e.
+a peak local emission rate $r_{n}^{\mathrm{wg,pk}}>\Gamma$ into the waveguide) can only occur for $n\geq\gamma/\Gamma+2$ [11], as we confirm in Fig.
+4a.
+Beyond this threshhold value, the asymptotic Dicke scaling $r_{n}^{\mathrm{wg,pk}}\sim n\Gamma$ is recovered for sufficiently large $n$.
+When $\tau>0$, we observe a more intricate interplay between collective and independent emission, owing to the additional loss of excitations into unguided modes during the propagation time of the photons in the waveguide.
+As shown in Fig.
+4a, at sufficiently small time delays the burst criterion from the retardation-free regime still holds and at sufficiently large $n$ the plateau effect can be observed.
+However, the plateau emission rate $r^{\mathrm{wg}}_{\mathrm{eff}}$ varies with $\gamma$ and, most notably, there exists a critical free-space decay rate $\gamma_{\mathrm{max}}$ that depends on the retardation time $\tau$, such that for larger $\gamma>\gamma_{\mathrm{max}}$, the burst vanishes entirely for all values of $n$ and $r_{n}^{\mathrm{wg,pk}}=\Gamma$ (see Fig.
+4b).
+It is worth noting that the emission rates $r_{n}^{\mathrm{wg}}(t)$ into the waveguide are not typically directly measurable in an experimental context.
+The local field intensity $I_{n}(t)$ in the waveguide (defined as in Eq.
+(7) of the main text, but with $r_{n}\to r_{n}^{\mathrm{wg}}$), represents a more experimentally accessible quantity.
+In Fig.
+4c, we show that the qualitative signatures of the plateau effect in the peak intensity $I_{n}^{\mathrm{pk}}$ are robust even in the presence of comparatively strong free-space decay.
+Figure 4: (a) Peak emission rate into the waveguide for $\tau=0$ (top) and $\Gamma\tau=5\times 10^{-4}$ (bottom).
+The retardation-free superradiance criterion from Ref.
+[11] is indicated in white.
+(b) Plateau emission rate $r_{\mathrm{eff}}^{\mathrm{wg}}$ as a function of free-space decay rate $\gamma$ for various time delays.
+(c) Local field intensity along the waveguide for $\Gamma\tau=10^{-3}$, both without free-space decay and with $\gamma/\Gamma\in\{1,2,5,10\}$, as indicated by the arrow.